OpenStudy (anonymous):

If p is a prime. Prove 12 is a factor of p^2-1, p>=5.

OpenStudy (anonymous):

fermat's little theorem with p = 2

OpenStudy (anonymous):

2 isn't greater than or equal to 5

OpenStudy (anonymous):

oh you are right this requires some thought. but i am fairly certain it is a direct application of fermat

OpenStudy (anonymous):

Probably

OpenStudy (anonymous):

but you are right, not trivial one. requires thinking

OpenStudy (anonymous):

oh no might not, might be easier. \[p^2-1=(p-1)(p+1)\] and p is odd so both factors are even and so the product is divisible by 4

OpenStudy (anonymous):

now we have to show it is divisible by 3 and then we are done right?

OpenStudy (anonymous):

and of course p is not divisible by 3 (it is prime) so \[p=3k+1\] or \[p=3k+2\] and now you can show that no matter which, \[p^2-1\] is a multiple of 3

OpenStudy (anonymous):

for example, if \[p=3k+1\] then \[(p+1)(p-1)=3k(p+1)\] hence divisible by 3

OpenStudy (anonymous):

last one is similar. let me know if this is not clear

OpenStudy (anonymous):

Nice!

OpenStudy (anonymous):

merci

OpenStudy (anonymous):

Well played. :)